# Weyl law for 1-cycles

### v1

@@ 4.2. Two important propositions to go from disks to manifolds
for every $k$-face $E^{\prime}$ of $E$. The same holds for $F|_{E}$ and together with the fact that $\partial B_{r_{E}}(e)\cap\mathcal{S}=\emptyset$ yields
$F(x)\llcorner\partial B_{r_{E}}(e)=0$
for every $x\in E$. We will extend $F^{\prime}|_{\partial E}$ to $E$ so that properties (D1) to (D5) hold. We will identify $E$ with the cone over $\partial E$ and we will use coordinates $(x,t)\in\partial E\times[0,1]$ where $(x,0)$ corresponds to $x$ and $(x,1)$ to the center of $E$ for each $x\in\partial E$. We define
$F^{\prime}(x,t)=F(x,t)\llcorner(\mathbb{D}^{3}\setminus B_{E})+F_{E}(x,t)$
where $B_{E}:=B_{r_{E}}(e)\cap\mathbb{D}^{3}$ and $F_{E}$ is supported in $B_{E}$ with $F_{E}(x,0)=F^{\prime}(x)\llcorner B_{E}$ for all $x\in\partial E$ so that $F^{\prime}(x,t)$ extends $F^{\prime}|_{\partial E}$. Thus our goal is to homotop $F^{\prime}(x)\llcorner B_{E}$ to a constant map while keeping properties (D1) to (D5).
Denote $S_{E}=\mathbb{S}^{2}\cap B_{E}$. As the number of points in $F_{E}(x,0)\llcorner(S_{E}\setminus B)$ is a priori unbounded for $x\in\partial E$, we can not just apply a homotopy $H:B_{E}\times[0,1]\to B_{E}$ contracting $B_{E}$ to the point $e$ to the family $F_{E}^{0}=F_{E}(x,0)$ because that could introduce too many points to $B\cup\operatorname{int}(\mathbb{D}^{3})$. Instead, we will contract $F_{E}^{0}$ in three steps.
@@
    The first one will homotop it to a family $F_{E}^{1}:\partial E\to\mathcal{Z}_{0}(\mathbb{D}^{3})$ supported in $\overline{B}_{L}(e)$ such that $F_{E}^{1}(x)\llcorner\partial B$ is supported in a discrete set $\{y_{1},...,y_{N}\}$ for every $x\in\partial E$ (the discrete set is independent of $x\in\partial E$). This homotopy will keep the property that
    $\operatorname{\mathbf{M}}(F_{E}^{1}(x,t)\llcorner\operatorname{int}(\mathbb{D}^{3}))+\operatorname{\mathbf{M}}(F_{E}^{1}(x,t)\llcorner\operatorname{int}(B))\leq M_{E}+k+1$
    where $F_{E}^{1}(x,0)=F_{E}^{0}(x)$ and $F_{E}^{1}(x,1)=F_{E}^{1}(x)$ for $x\in\partial C$.
@@
    The second one will homotop $F_{E}^{1}$ to a family $F_{E}^{2}$ also supported in $\overline{B}_{L}(e)$ such that $F_{E}^{2}(x)\llcorner\partial B$ is supported in a set $\{z\}$ of $1$ element. This homotopy will verify
    $\operatorname{\mathbf{M}}(F_{E}^{2}(x,t)\llcorner\operatorname{int}(\mathbb{D}^{3}))+\operatorname{\mathbf{M}}(F_{E}^{2}(x,t)\llcorner\operatorname{int}(B))\leq M_{E}+k+2.$
@@
    The third homotopy will just contract $F_{E}^{2}$ radially towards $e$ in $\overline{B}_{L}(e)$ (mapping points of $\partial\mathbb{D}^{3}$ into $\partial\mathbb{D}^{3}$), increasing the mass bound from $M_{E}+k+1$ to at most $M_{E}+k+2$.
@@
We proceed to describe how to construct the first two homotopies, being them induced by deformation retractions of subspaces of $B_{E}$. For (1), let $T_{1},...,T_{N}$ be an enumeration of $\{T_{C}:C\in\operatorname{Faces}_{p}(X),C\supseteq E\}$ according to an increasing ordering of the corresponding $\theta_{C}$. Let $T^{\prime}_{1},...,T^{\prime}_{N}$ be intermediate $T_{\theta,\eta}$ with $\theta^{\prime}_{i}\in(\theta_{i},\theta_{i+1})$ modulo $2\pi$ and $\eta\leq\eta_{i}$ for every $1\leq i\leq N$. By inductive hypothesis, $F_{E}^{0}(x)=F^{\prime}(x)\llcorner B_{E}$ is supported in $B_{E}^{*}:=B_{E}\setminus\cup_{i=1}^{N}T_{i}$ for $x\in\partial E$. Given $0<r_{1}<r_{2}<\frac{\pi}{2}$, let
$\Omega_{r_{1},r_{2}}=\{(x,y,z)\in\mathbb{D}^{3}:-\cos(r_{1})\leq z\leq-\cos(r_{2})\}$
and
$\Omega_{r_{1}}=\{(x,y,z)\in\mathbb{D}^{3}:z\leq-\cos(r_{1})\}.$
@@
Consider a homotopy $H_{E}^{1}:B_{E}^{*}\times[0,1]\to B_{E}^{*}$ which is the identity in $\Omega_{\frac{L}{2}}$ and acts as follows in $\Omega_{L,r_{E}}$. Each $z$-slice of $\Omega_{L,r_{E}}$ is a circle $C_{r}$ with a certain radius $r$. The homotopy $H_{E}^{1}$ restricted to each of those slices retracts $C_{r}\setminus\cup_{i=1}^{N}T_{i}$ (which is a circle minus $N$ straight short segments emanating from boundary points) to a smaller concentric circle $C_{r}^{\prime}$ union $N$ segments $T^{\prime}_{i}\cap C_{r}$. The homotopy is chosen so that points in $\partial C_{r}$ remain in $\partial C_{r}$ and at the end are mapped to $\partial C_{r}\cap\cup_{i=1}^{N}T^{\prime}_{i}$. For $-\cos(\frac{L}{2})\leq z\leq-\cos(L)$, we interpolate between the $H_{E}^{1}$ previously defined and the identity. Observe that the points in $H_{E}(F_{E}^{0}(x),1)\llcorner(S_{E}\setminus B)$ lie in the $N$ geodesic segments $T_{i}^{\prime}\cap\mathbb{S}^{2}$. Then our next step is to retract $C_{L,r_{E}}$ onto $\partial B$ radially towards $e$. Hence we obtain a family $F_{E}^{1}$ which is supported in $\Omega_{L}$ with the property that $F_{E}^{1}\llcorner\partial B$ is supported in $\{y_{1},...,y_{N}\}$ where $y_{i}=T_{i}^{\prime}\cap\partial B$. A more rigorous description of these homotopies will be provided when extending the result to dimension $n$.
@@
To obtain $F_{E}^{2}(x,t)$, we need to apply $N-1$ homotopies $\overline{H}_{E}^{1},...,\overline{H}_{E}^{N-1}$ to $F_{E}^{1}$ so that if $\overline{F}_{E}^{i}$ is the map obtained after applying the first $i$, $\overline{F}_{E}^{i}(x)\llcorner(\partial{B}_{L}(e)\cap\mathbb{S}^{2})$ is supported in $\{y_{i+1},...,y_{N}\}$. We describe $\overline{H}_{E}^{i}$ in the circle $C=\mathbb{D}^{3}\cap\{z=-\cos(L)\}$. We consider a curve $\gamma$ connecting $y_{i}$ with $y_{i+1}$ in $C$ without touching any of the $T_{i}$. We take a tiny “tubular neighborhood” $N_{\gamma}$ of $\gamma$ in $C$ which also avoids the $T_{i}$ and only intersects $\{y_{1},...,y_{N}\}$ at $\{y_{i},y_{i+1}\}$. It is possible to retract $C$ onto $C\setminus N_{\gamma}$ so that the trajectory of $y_{i}$ is given by $\gamma$. Then we can extend this to a homotopy $\overline{H}_{E}^{i}$ in $B_{L}(e)$ which is not the identity only in a tiny neighborhood $\tilde{N}_{\gamma}\supseteq N_{\gamma}$ of $\gamma$ in $B_{L}(e)$ (and in particular, away from the $T_{i}$). We can think of the deformation induced by $\overline{H}_{E}^{i}$ as that of digging a narrow tunnel in $B_{L}(e)$ along the curve $\gamma$ starting at the point $y_{i}$ and ending at $y_{i+1}$. By applying to the family $F_{2}^{E}$ the composition of the $H_{E}^{i}$ we obtain a map with the properties in (2) above.
@@
Let $\mathbb{S}^{n-1}\subseteq\mathbb{R}^{n}$ be the unit sphere. Let $P^{d}\subseteq\mathbb{S}^{n-1}$, $11\leq d\leq n-$ 1, be an embedded $d$-dimensional polyhedron in $\mathbb{S}^{n-1}$ whose cells are geodesically convex in $\mathbb{S}^{n-1}$. Suppose that a certain hyperplane $\Pi\subseteq\mathbb{R}^{n}$ cuts $P$. Then $\Pi$ cuts the $k$-skeleton of $P$ for every $1\leq k\leq d$.
@@
We proceed by induction in $d$. If $d=1$ it is automatic. Suppose it is true for $d-1$ and consider a $d$-dimensional polyhedron $P\subseteq\mathbb{S}^{n-1}$ and a hyperplane $\Pi\subseteq\mathbb{R}^{n}$ which intersects $P$. We want to show first that $\Pi\cap P_{d-1}\neq\emptyset$. Suppose the contrary. Pick $C$ a top dimensional cell of $P$ such that $\Pi\cap C\neq\emptyset$ and a point $p\in\Pi\cap C$. It must be $p\in\operatorname{int}(C)$ as $\partial C\subseteq P_{d-1}$. Denote $H^{+},H^{-}$ the two open half spaces determined by $\Pi$. We have
$\partial C=\partial C\cap H^{+}\cup\partial C\cap H^{-}$
as a disjoint union of open sets, so by connectedness it must be $\partial C\subseteq H^{+}$ or $\partial C\subseteq H^{-}$. Assume the first is true without loss of generality. Then as $H^{+}\cap\mathbb{S}^{n-1}$ is geodesically convex and contains the vertices of the cell $C$, it must contain all of $C$. But this is absurd because $p\notin H^{+}$ (as it is in $\Pi$). Therefore, $\Pi\cap P_{d-1}\neq\emptyset$ and by inductive hypothesis $\Pi$ intersects $P_{k}$ for every $11\leq k\leq d-$ 1 as desired. ∎
@@
For each $i\in I$, let $Q_{i}$ be an embedded convex $(n-1)$-cube in $\mathbb{S}^{n-1}$ containing $B_{i}$ (being $Q_{i}$ the convex hull of its $2^{n-1}$ vertices) whose diameter is at most $d(n)\operatorname{diam}(B_{i})$ for a certain dimensional constant $d(n)$. Let $Q=\bigcup_{i\in I}Q_{i}$. Observe that if a hyperplane $\Pi\subseteq\mathbb{R}^{n}$ intersects $\bigcup_{i\in I}B_{i}$ then it intersects $\bigcup_{i\in I}Q_{i}$ and hence by Lemma 4.9 (applied to the $Q_{i}$) it also cuts its $1$-skeleton $Q_{1}$. Therefore it suffices to find a hyperplane $\Pi\subseteq\mathbb{R}^{n}$ which does not intersect $Q_{1}$. Denote $n(\Pi,Q_{1})$ the number of points in $\Pi\cap Q_{1}$. By Crofton’s formula, if $\mu$ denotes the rotation-invariant metric on the space $\mathcal{P}$ of planes through the origin then
[…]
Therefore, if every hyperplane $\Pi\in\mathcal{P}$ intersected $Q_{1}$, as $n(\Pi,Q_{1})$ is integer valued we would have
[…]
and hence $\operatorname{L}(Q_{1})\geq\frac{\mu(\mathcal{P})}{c(n)}$. But
$\operatorname{L}(Q_{1})=\sum_{i\in I}\operatorname{L}((Q_{i})_{1})\leq\sum_{i\in I}e(n)\operatorname{diam}(Q_{i})\leq e(n)d(n)\sum_{i\in I}\operatorname{diam}(B_{i})$
where $e(n)$ denotes the number of $1$-faces of an $n$-dimensional cube. Thus taking $\delta_{n}=\frac{\mu(\mathcal{P})}{c(n)d(n)e(n)}$ we get the desired result. ∎
@@
Let $F:X^{p}\to\mathcal{Z}_{0}(\mathbb{D}^{n})$ be a $\delta$-localized family. Let $L>0$ be a very small real number and let $B$ be the open ball in $\mathbb{S}^{n-1}$ centered at $e$ of radius $L$. There exists a continuous family $F^{\prime}:X^{p}\to\mathcal{Z}_{0}(\mathbb{D}^{n})$ with the following properties:
@@
    If $\tilde{B}=B_{c(n,p)\delta}(e)$ then $F^{\prime}(x)\llcorner(\mathbb{D}^{n}\setminus\tilde{B})=F(x)\llcorner(\mathbb{D}^{n}\setminus\tilde{B})$ for every $x\in X^{p}$.
@@
Given $l\in[0,\frac{\pi}{2})$ denote
$\mathbb{S}^{n-1}_{l}=\{x\in\mathbb{S}^{n-1}:-\cos(l)\leq x_{n}\leq 0\}$
and
$\mathbb{S}^{n-2}_{l}$ =$=\{x\in\mathbb{S}^{n-1}:x_{n}=-\cos(l)\}$
$=\mathbb{S}^{n-1}\cap\partial B_{l}(e)$
$=\partial\mathbb{S}^{n-1}_{l}$
where $B_{l}(e)$ denotes the ball in $\mathbb{S}^{n-1}$ of radius $l$ centered at $e$. Given $\varepsilon\in(0,\frac{1}{2})$ we define
$R(l,\varepsilon)=\{(1-s)x:x\in\mathbb{S}^{n-1}_{l},s\in[0,\varepsilon]\}.$
@@
For every $0<l<\frac{\pi}{2}$ the map $\phi:\mathbb{S}^{n-2}\times[-\cos(l),0]\to\mathbb{S}^{n-1}_{l}$ given by
@@
is a diffeomorphism. Additionally, for each $\varepsilon\in(0,\frac{1}{2})$ we can extend the previous to a map $\phi:\mathbb{S}^{n-2}\times[-\cos(l),0]\times[0,\varepsilon]\to R(l,\varepsilon)$ (regarding $\mathbb{S}^{n-2}\times[-\cos(l),0]\subseteq\mathbb{S}^{n-2}\times[-\cos(l),0]\times[0,\varepsilon]$ via the embedding $(u,t)\mapsto(u,t,0)$) in the following way
@@
Hence we can use $(u,t,s)\in\mathbb{S}^{n-2}\times[-\cos(l),0]\times[0,\varepsilon]$ as a coordinate system for $R(l,\varepsilon)$. Notice that if $p=\phi(u,t,s)$ the coordinate $s$ measures the distance from $p$ to $\mathbb{S}^{n-1}$ and if $s=0$, $t$ is the $x_{n}$-coordinate of $p$ and $u$ represents its location in the parallel $\mathbb{S}^{n-1}\cap\{x_{n}=t\}$.
@@
Notice that when $n=3$, $S_{\Pi}$ is determined by two antipodal points in $\mathbb{S}^{1}$ and hence it can be associated with a certain angle $\theta\in[0,\pi)$; being somehow a generalization of the $s_{\theta}$ defined before (which corresponded to a certain angle $\theta\in[0,2\pi)$). This also yields an analogy between $T_{\Pi,\varepsilon}$ and $T_{\theta,\varepsilon}$.
@@
Observe that $S_{\Pi}\subseteq\Pi\times\mathbb{R}\cap\mathbb{S}^{n-1}$ and $T_{\Pi,\varepsilon}\subseteq\Pi\times\mathbb{R}\cap\mathbb{D}^{n}$, moreover
$S_{\Pi,\varepsilon}$ =$=\{x\in\Pi\times\mathbb{R}\cap\mathbb{S}^{n-1}:x_{n}\in[-\cos(\frac{L}{2}),0]\}$
$T_{\Pi,\varepsilon}$ =$=\{x=\phi(u,t,s)\in\Pi\times\mathbb{R}\cap\mathbb{D}^{n}:t\in[-\cos(\frac{L}{2}),0],s\in[0,\varepsilon]\}\}.$
Definition 4.17.
In this section, for each $r<\frac{\pi}{2}$ we denote
$B_{r}(e)=\{sw:w\in\mathbb{S}^{n-1},\operatorname{dist}_{\mathbb{S}^{n-1}}(w,e)<r,s\in(1-r,1]\}.$

### v2

@@ 4.2. Two important propositions to go from disks to manifolds
for every $k$-face $E^{\prime}$ of $E$. The same holds for $F|_{E}$ and together with the fact that $C_{E}\cap\mathcal{S}=\emptyset$ yields
$F(x)\llcorner C_{E}=0$
for every $x\in E$. In particular, $F(x)\llcorner(\mathbb{D}^{3}\setminus U_{E})$ is a continuous function on $E$, and the same holds for $F^{\prime}(x)\llcorner U_{E}$ on $\partial E$. We will extend $F^{\prime}|_{\partial E}$ to $E$ so that properties (F1) to (F5) hold. We will identify $E$ with the cone over $\partial E$ and we will use coordinates $(x,t)\in\partial E\times[0,1]$ where $(x,0)$ corresponds to $x$ and $(x,1)$ to the center of $E$ for each $x\in\partial E$. We define
$F^{\prime}(x,t)=F(x,t)\llcorner(\mathbb{D}^{3}\setminus U_{E})+F_{E}(x,t)$
with $F_{E}$ supported in $U_{E}$ and $F_{E}(x,0)=F^{\prime}(x)\llcorner U_{E}$ for all $x\in\partial E$. The previous will imply that $F^{\prime}(x,t)$ extends $F^{\prime}|_{\partial E}$. Thus, our goal is to homotop $F_{E}^{0}:\partial E\to\mathcal{Z}_{0}(U_{E})$ given by $F_{E}^{0}(x)=F^{\prime}(x)\llcorner U_{E}$ to a constant map while keeping properties (F1) to (F5).
Denote $S^{1}_{E}=\partial B_{E}$. As the number of points in $F_{E}^{0}(x)\llcorner(B_{E}\setminus B)$ is a priori unbounded for $x\in\partial E$, as explained before we can not just apply a homotopy $H_{E}:U_{E}\times[0,1]\to U_{E}$ contracting $U_{E}$ to the point $e$ to the family $F_{E}^{0}$ because that could introduce too many points to $B\cup\operatorname{int}(\mathbb{D}^{3})$. Instead, we will contract $F_{E}^{0}$ in three steps.
@@
    The first one will homotop it to a family $F_{E}^{1}:\partial E\to\mathcal{Z}_{0}(\mathbb{D}^{3})$ supported in $\overline{U}$ ($U=U_{L}$) such that $F_{E}^{1}(x)\llcorner\partial B$ is supported in a discrete set $\{y_{1},...,y_{N}\}$ for every $x\in\partial E$ (the discrete set is independent of $x\in\partial E$). The mentioned homotopy $F_{E}^{0}:\partial E\times[0,1]\to\mathcal{Z}_{0}(B_{E})$ will keep the property that
    $\operatorname{\mathbf{M}}(F_{E}^{0}(x,t)\llcorner\operatorname{int}(\mathbb{D}^{3}))+\operatorname{\mathbf{M}}(F_{E}^{0}(x,t)\llcorner\operatorname{int}(B))\leq M_{E}+k+1$
    where $F_{E}^{0}(x,0)=F_{E}^{0}(x)$ and $F_{E}^{0}(x,1)=F_{E}^{1}(x)$ for $x\in\partial E$.
@@
    The second one will homotop $F_{E}^{1}$ to a family $F_{E}^{2}$ also supported in $\overline{U}$ such that $F_{E}^{2}(x)\llcorner\partial B$ is supported in a set $\{q\}$ of $1$ element. The corresponding homotopy $F_{E}^{1}:\partial E\times[0,1]\to\overline{U}$ will verify
    $\operatorname{\mathbf{M}}(F_{E}^{1}(x,t)\llcorner\operatorname{int}(\mathbb{D}^{3}))+\operatorname{\mathbf{M}}(F_{E}^{1}(x,t)\llcorner\operatorname{int}(B))\leq M_{E}+k+2.$
@@
    The third homotopy will just contract $F_{E}^{2}$ radially towards $e$ in $\overline{U}$ (mapping points of $\mathbb{S}^{2}$ into $\mathbb{S}^{2}$), increasing the mass bound from $M_{E}+k+1$ to at most $M_{E}+k+2$.
@@
We proceed to describe how to construct the first two homotopies, being them induced by deformation retractions of subspaces of $U_{E}$. For (1), let $T_{1},...,T_{N}$ be an enumeration of $\{T_{C}:C\in\operatorname{Faces}_{p}(X),C\supseteq E\}$ according to an increasing ordering of the corresponding $\theta_{C}$. Let $T^{\prime}_{1},...,T^{\prime}_{N}$ be intermediate $T_{\theta,\eta}$ with $\theta^{\prime}_{i}\in(\theta_{i},\theta_{i+1})$ modulo $2\pi$ and $\eta\leq\eta_{i}$ for every $1\leq i\leq N$. By inductive hypothesis, $F_{E}^{0}(x)=F^{\prime}(x)\llcorner U_{E}$ is supported in $U_{E}^{*}:=U_{E}\setminus\cup_{i=1}^{N}T_{i}$ for $x\in\partial E$. Consider a homotopy $H_{E}^{1}:U_{E}^{*}\times[0,1]\to U_{E}^{*}$ which is the identity in $U_{\frac{L}{2}}$ and acts as follows in $U_{r_{E}}\setminus U_{\frac{L}{2}}$. We can write
$U_{r_{E}}\setminus U_{\frac{L}{2}}=\bigcup_{r\in[\frac{L}{2},r_{E}]}C_{r}$
@@
as a disjoint union (each $C_{r}$ corresponds to a $z$-slice of $U_{r_{E}}\setminus U_{\frac{L}{2}}$). For $r\in[L,r_{E}]$, the homotopy $H_{E}^{1}$ restricted to $C_{r}$ retracts $C_{r}\setminus\cup_{i=1}^{N}T_{i}$ (which is a circle minus $N$ straight short segments emanating from boundary points) to the smaller concentric circle $C_{r}^{\prime}$ of radius $\operatorname{rad}(C_{r})-\eta$ union the $N$ segments $T^{\prime}_{i}\cap C_{r}$ (each of them of length $\eta$). The homotopy is chosen so that points in $\partial C_{r}$ remain in $\partial C_{r}$ and at the end they are mapped to $\partial C_{r}\cap\cup_{i=1}^{N}T^{\prime}_{i}$ (which is a finite set of points). For $r\in[\frac{L}{2},L]$, we interpolate between the $H_{E}^{1}$ previously defined and the identity. Observe that the points in $H_{E}^{1}(F_{E}^{0}(x),1)\llcorner(B_{E}\setminus B)$ lie in the $N$ geodesic segments $T_{i}^{\prime}\cap\mathbb{S}^{2}$. Then our next step is to retract $U_{E}$ onto $U$ radially towards $e$. Hence, we obtain a family $F_{E}^{1}$ which is supported in $\overline{U}$ with the property that $F_{E}^{1}\llcorner\partial B$ is supported in $\{y_{1},...,y_{N}\}$ where $y_{i}=T_{i}^{\prime}\cap\partial B$. A more rigorous description of these homotopies will be provided when the result is extended to $n\geq 4$.
@@
To obtain $F_{E}^{2}(x,t)$, we need to apply N1N-1 homotopies $\overline{H}_{E}^{1},...,\overline{H}_{E}^{N-1}$ to $F_{E}^{1}$ so that if $\overline{F}_{E}^{i}$ is the map obtained after applying the first $i$, $\overline{F}_{E}^{i}(x)\llcorner\partial B$ is supported in $\{y_{i+1},...,y_{N}\}$. We describe $\overline{H}_{E}^{i}$ in the circle $C_{L}$. We consider a curve $\gamma$ connecting $y_{i}$ with $y_{i+1}$ in $C_{L}$ without touching any of the $T_{i}$. We take a tiny “tubular neighborhood” $N_{\gamma}$ of $\gamma$ in $C_{L}$ which also avoids the $T_{i}$ and only intersects $\{y_{1},...,y_{N}\}$ at $\{y_{i},y_{i+1}\}$. It is possible to retract $C_{L}$ onto $C_{L}\setminus N_{\gamma}$ so that the trajectory of $y_{i}$ is given by $\gamma$. Then we can extend this to a homotopy $\overline{H}_{E}^{i}$ in $\overline{U}$ which is not the identity only in a tiny neighborhood $\tilde{N}_{\gamma}\supseteq N_{\gamma}$ of $\gamma$ in $\overline{U}$ (and in particular, away from the $T_{i}$). We can think of the deformation induced by $\overline{H}_{E}^{i}$ as that of digging a narrow tunnel in $U$ along the curve $\gamma$ starting at the point $y_{i}$ and ending at $y_{i+1}$. By applying to the family $F_{2}^{E}$ the composition of the $\overline{H}_{E}^{i}$ we obtain a map with the properties in (2) above.
@@
Let $\mathbb{S}^{n-1}\subseteq\mathbb{R}^{n}$ be the unit sphere. Let $P^{d}\subseteq\mathbb{S}^{n-1}$, $1\leq d\leq n-1$, be an embedded $d$-dimensional polyhedron in $\mathbb{S}^{n-1}$ whose cells are geodesically convex in $\mathbb{S}^{n-1}$. Suppose that a certain hyperplane $\Pi\subseteq\mathbb{R}^{n}$ intersects $P$. Then $\Pi$ intersects the $k$-skeleton of $P$ for every $1\leq k\leq d$.
@@
We proceed by induction in $d$. If $d=1$ it is automatic. Suppose it is true for d1d-1 and consider a $d$-dimensional polyhedron $P\subseteq\mathbb{S}^{n-1}$ and a hyperplane $\Pi\subseteq\mathbb{R}^{n}$ which intersects $P$. We want to show first that $\Pi\cap P_{d-1}\neq\emptyset$. Suppose the contrary. Pick $Q$ a top dimensional cell of $P$ such that $\Pi\cap Q\neq\emptyset$ and a point $q\in\Pi\cap Q$. It must be $q\in\operatorname{int}(Q)$ as $\partial Q\subseteq P_{d-1}$. Denote $H^{+},H^{-}$ the two open half spaces of $\mathbb{R}^{n}$ determined by $\Pi$. We have
$\partial Q=\partial Q\cap H^{+}\cup\partial Q\cap H^{-}$
as a disjoint union of open sets, so by connectedness it must be $\partial Q\subseteq H^{+}$ or $\partial C\subseteq H^{-}$. Assume the first is true without loss of generality. Then as $H^{+}\cap\mathbb{S}^{n-1}$ is geodesically convex and contains the vertices of the cell $Q$, it must contain all of $Q$. But this is absurd because $q\notin H^{+}$ (as it is in $\Pi$). Therefore, $\Pi\cap P_{d-1}\neq\emptyset$ and by inductive hypothesis $\Pi$ intersects $P_{k}$ for every $1\leq k\leq d-1$ as desired. ∎
@@
For each $i\in I$, let $Q_{i}$ be an embedded convex (n1)(n-1)-simplex in $\mathbb{S}^{n-1}$ containing $B_{i}$ whose diameter is at most $d(n)\operatorname{diam}(B_{i})$ for a certain dimensional constant $d(n)$. Let $Q_{(1)}=\bigcup_{i\in I}(Q_{i})_{1}$ be the union of the $1$-skeletons of the $Q_{i}$. Observe that if a hyperplane $\Pi\subseteq\mathbb{R}^{n}$ intersects $\bigcup_{i\in I}B_{i}$ then it intersects $Q_{i}$ for some $i\in I$ and hence by Lemma 4.9 it also cuts its $1$-skeleton, which implies that it intersects $Q_{(1)}$. Therefore, it suffices to find a hyperplane $\Pi\subseteq\mathbb{R}^{n}$ which does not intersect $Q_{(1)}$. Denote $n(\Pi,Q_{(1)})$ the number of points in $\Pi\cap Q_{(1)}$. By Crofton’s formula, there is a universal constant $C(n)>0$ such that if $\mu$ denotes the rotation-invariant metric on the space $\mathcal{P}$ of planes through the origin then
$\int_{\mathcal{P}}n(\Pi,Q_{(1)})d\mu(\Pi)=C(n)\operatorname{L}(Q_{(1)}).$
Therefore, if every hyperplane $\Pi\in\mathcal{P}$ intersected $Q_{(1)}$, as $n(\Pi,Q_{(1)})$ is integer valued we would have
$\mu(\mathcal{P})=\int_{\mathcal{P}}1d\mu(\Pi)\leq\int_{\mathcal{P}}n(\Pi,Q_{(1)})d\mu(\Pi)=C(n)\operatorname{L}(Q_{(1)})$
and hence $\operatorname{L}(Q_{(1)})\geq\frac{\mu(\mathcal{P})}{C(n)}$. But
$\operatorname{L}(Q_{(1)})=\sum_{i\in I}\operatorname{L}((Q_{i})_{1})\leq\sum_{i\in I}e(n)\operatorname{diam}(Q_{i})\leq e(n)d(n)\sum_{i\in I}\operatorname{diam}(B_{i})$
where $e(n)={n\choose 2}$ denotes the number of $1$-faces of an $n$-dimensional simplex. Thus taking $\delta_{n}=\frac{\mu(\mathcal{P})}{C(n)d(n)e(n)}$ we get the desired result. ∎
@@
There exists $\delta_{0}=\delta_{0}(n,p)>0$ such that the following is true. Let $F:X^{p}\to\mathcal{Z}_{0}(\mathbb{D}^{n})$ be a $\delta$-localized family, $\delta<\delta_{0}$. Let $0<L<\delta$ be a very small real number and let $B$ be the open ball in $\mathbb{S}^{n-1}$ centered at $e$ of radius $L$. There exists a continuous family $F^{\prime}:X^{p}\to\mathcal{Z}_{0}(\mathbb{D}^{n})$ with the following properties:
@@
[…]
@@
Given $r\in[0,\frac{\pi}{2})$ and $\varepsilon\in(0,\frac{1}{2})$, denote
$B_{r}$ $=B^{\mathbb{S}^{n-1}}_{r}(e),$
$U_{r}$ $=\{(1-s)q:q\in B_{r},s\in[0,r]\},$
$S^{n-2}_{r}$ $=\{x\in\mathbb{S}^{n-1}:x_{n}=-\cos(r)\},$
$S^{n-1}_{r}$ $=\{(1-s)q:q\in S^{n-2}_{r},s\in[0,r]\}\cup\{(1-r)q:q\in B_{r}\},$
$A^{n-1}_{r}$ $=\{x\in\mathbb{S}^{n-1}:-\cos(r)\leq x_{n}\leq 0\},$
$A^{n}_{r,\varepsilon}$ $=\{(1-s)q:q\in A^{n-1}_{r},s\in[0,\varepsilon]\}$
where $B_{r}^{\mathbb{S}^{n-1}}(e)$ denotes the ball in $\mathbb{S}^{n-1}$ of radius $r$ centered at $e$.
@@
Observe that
$\partial B_{r}$ $=S^{n-2}_{r},$
$\partial U_{r}$ $=S^{n-1}_{r}\cup B_{r},$
$A^{n-1}_{r}$ $=A^{n}_{r,\varepsilon}\cap\mathbb{S}^{n-1}.$
In addition, $B=B_{L}$ and we denote $U=U_{L}$.
Remark 4.14.
For every $0<r<\frac{\pi}{2}$ the map $\phi:\mathbb{S}^{n-2}\times[-\cos(r),0]\to A^{n-1}_{r}$ given by
@@
is a diffeomorphism. Furthermore, for each $\varepsilon\in(0,\frac{1}{2})$ we can extend the previous to a map $\phi:\mathbb{S}^{n-2}\times[-\cos(r),0]\times[0,\varepsilon]\to A^{n}_{r,\varepsilon}$ (regarding $\mathbb{S}^{n-2}\times[-\cos(r),0]\subseteq\mathbb{S}^{n-2}\times[-\cos(r),0]\times[0,\varepsilon]$ via the embedding $(u,t)\mapsto(u,t,0)$) in the following way
@@
Hence we can use $(u,t,s)\in\mathbb{S}^{n-2}\times[-\cos(r),0]\times[0,\varepsilon]$ as a coordinate system for $A^{n}_{r,\varepsilon}$. Notice that if $q=\phi(u,t,s)$ the coordinate $s$ measures the distance from $q$ to $\mathbb{S}^{n-1}$ and if $s=0$, $t$ is the $x_{n}$-coordinate of $q$ and $u$ represents its location in the parallel $\mathbb{S}^{n-1}\cap\{x_{n}=t\}$.
@@
Notice that when $n=3$, $s_{\Pi}$ is determined by two antipodal points in $\mathbb{S}^{1}$ and hence can be associated with a certain angle $\theta\in[0,\pi)$; being a generalization of the $s_{\theta}$ defined before (which corresponded to a certain angle $\theta\in[0,2\pi)$). This also yields an analogy between $T_{\Pi,\varepsilon}$ and $T_{\theta,\varepsilon}$.
@@
Observe that $s_{\Pi}\subseteq(\Pi\times\mathbb{R})\cap\mathbb{S}^{n-1}$ and $T_{\Pi,\varepsilon}\subseteq(\Pi\times\mathbb{R})\cap\mathbb{D}^{n}$, moreover
$s_{\Pi}$ $=\{x\in(\Pi\times\mathbb{R})\cap\mathbb{S}^{n-1}:x_{n}\in[-\cos(\frac{L}{2}),0]\}$
$T_{\Pi,\varepsilon}$ $=\{x=\phi(u,t,s)\in(\Pi\times\mathbb{R})\cap\mathbb{D}^{n}:t\in[-\cos(\frac{L}{2}),0],s\in[0,\varepsilon]\}\}.$